# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### v2

@@ Acknowledgments
Appendix A Dependency of cavity-EIT linewidth on cavity coupling parameter $g$ and control field Rabi frequency, $\Omega_{c}$ for different $n_{th}$
@@
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 7 For a weakly excited motional state ($n_{th}=0.5$) (see 7(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 7(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 7(c), 7(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.
@@
As the control-field strength increases, it partially mitigates this effect by stabilizing the dark state against motional fluctuations. The strong sensitivity of the cavity response to the ions motional state is highlighted by the systematic and temperature-dependent modification of the cavity-EIT linewidth shown in these plots. This provide a solid foundation for thermometry using cavity-based EIT.
Appendix B Verification of different sideband transition related phenomenon
As a proof of principle of our approach, we performed verification of different sideband transition related characteristics. In case of sideband transitions, a two-level system dressed with vibrational levels is sufficient to address different properties. The sideband Hamiltonians used here are consistent with the Lamb–Dicke-expanded interaction employed in the main text. The two kinds of sideband transitions (BSB and RSB) are described by the Hamiltonian which is similar to the 4th term (with the multiplication of the Lamb-Dicke parameter $\eta$) of Equation 2 with the reduced Hilbert space of only photon and phonon. The Hamiltonian for blue and red sideband transition can be written as [1],
@@ Appendix B Verification of different sideband transition related phenomenon
We verified this principle for a range of phonon numbers and different Lamb-Dicke parameters. In Figure 8, the comparison between the probability ratio from simulation and phonon ratio from Equation 14 has been shown for four different Lamb-Dicke parameters. At lower calues of Lamb-Dicke parameters, these two quantities are well-matched, which consistent with the theory. For higher Lamb-Dicke parameters the simulation result and theory starts to deviate as the boundary of Lamb-Dicke regime ($\eta\ll 1$) is approached, whereas sideband transitions assume that the ion is well within the Lamb-Dicke Regime.
Figure 9: Rabi oscillation in blue sideband transition for $n=0\rightarrow 1$. The red dots represent the data from the NIST experiment [9], and the solid blue line shows the simulation result for the same set of parameters used in the experiment. The parameters are, $\eta=0.202$, $\Omega=2\pi\times 500$ kHz, $\gamma=20$ kHz.
The dephasing and frequency in the Rabi oscillation of the blue sideband transition depend on the phonon occupation. There are many experimental verifications for this phenomenon [5, 9]. We have verified this argument and matches with the experimental data reported in the NIST experiment (see Figure 9) [9].

### v3

@@ Appendix B Comparison of simulation results with analytical expressions and experimental data
Appendix C Dependency of cavity-EIT linewidth on cavity coupling parameter $g$ and control field Rabi frequency $\Omega_{c}$ for different $n_{th}$
@@
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 12 For a weakly excited motional state ($n_{th}=0.5$) (see 12(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 12(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 12(c), 12(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.
@@
As the control-field strength increases, it partially mitigates this effect by stabilizing the dark state against motional fluctuations. The strong sensitivity of the cavity response to the ions’ motional state is highlighted by the systematic and temperature-dependent modification of the cavity-EIT linewidth shown in these plots. This provides a solid foundation for thermometry using cavity-based EIT.
Appendix D Verification of different sideband transition -related phenomenon
As proof of principle of our approach, we performed verification of different sideband transition related characteristics. In the case of sideband transitions, a two-level system dressed with vibrational levels is sufficient to address different properties. The sideband Hamiltonians used here are consistent with the Lamb–Dicke-expanded interaction employed in the main text. The two kinds of sideband transitions (BSB and RSB) are described by the Hamiltonian which is similar to the 4th term (with the multiplication of the Lamb-Dicke parameter $\eta$) of Equation 10 with the reduced Hilbert space of only photon and phonon. The Hamiltonian for blue and red sideband transition can be written as [1],
@@ Appendix D Verification of different sideband transition -related phenomenon
We verified this principle for a range of phonon numbers and different Lamb-Dicke parameters. In Figure 13, the comparison between the probability ratio from simulation and phonon ratio from Equation 22 has been shown for four different Lamb-Dicke parameters. At lower values of Lamb-Dicke parameters, these two quantities are well-matched, which consistent with the theory. For higher Lamb-Dicke parameters the simulation result and theory start to deviate as the boundary of the Lamb-Dicke regime ($\eta\ll 1$) is approached, whereas sideband transitions assume that the ion is well within the Lamb-Dicke regime.
Figure 14: Rabi oscillation in blue sideband transition for $n=0\rightarrow 1$. The red dots represent the data from the NIST experiment [9], and the solid blue curve represents the simulated result obtained using the same parameters as those employed in the experiment. The parameters are, $\eta=0.202$, $\Omega=2\pi\times 500$ kHz, $\gamma=20$ kHz.
The dephasing and frequency in the Rabi oscillation of the blue sideband transition depend on the phonon occupation. There are many experimental verifications for this phenomenon [7, 9]. We have verified this argument, and it matches the experimental data reported in the NIST experiment (see Figure 14) [9].